OpenStudy (anonymous):

Billie will receive one-sixth of a $1000 reward for finding a stolen purse as well as an additional one-fourth of the reward for identifying the thief. What is the total portion of the reward which Billie will receive?

OpenStudy (anonymous):

1000 * 1/6 = 166.67 1000 * 1/4 = 250 166.67 + 250 = 416.67

OpenStudy (anonymous):

how did you get the 166.67?

OpenStudy (anonymous):

1000/1 * 1/6 = 1000/6...So, basically....1000 divided by 6

OpenStudy (anonymous):

(1/6)*1000 = one sixth of one thousand. If we evaluate that, it comes to 1000/6, both numerator and denominator can be divided by 2 for reduction: 500/3. Since a third of anything is equal to .33333333 to infinity, the more correct answer would be to leave it at 500/3.

OpenStudy (anonymous):

you got one sixth of 1000

OpenStudy (anonymous):

However, because they asked us to get dollars: 500/3 is equal to 500 / 3 = 166.67 dollars. Hope that helps!

OpenStudy (anonymous):

i this problem do I need to reduced it?

OpenStudy (anonymous):

You only need to reduce if your answer winds up in a fraction form and is reduceable, which this one doesn't....This one is in decimal form...So, you're okay. :D

OpenStudy (anonymous):

thanks